40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue May 29 14:58:28 2012, Jenne, Update, Locking, ETMX trans camera 
    Reply  Tue May 29 15:49:02 2012, Koji, Update, Locking, ETMX trans camera 
       Reply  Thu May 31 11:02:14 2012, steve, Update, Locking, ETMX 1064 trans camera 
          Reply  Wed Jun 20 16:37:11 2012, yuta, Update, Locking, ETMX 1064 trans camera 
             Reply  Thu Jun 21 12:56:49 2012, yuta, Update, Locking, ETMX 1064 trans camera Xarm_EndTableLayout_NewTransCamera.png
Message ID: 6838     Entry time: Wed Jun 20 16:37:11 2012     In reply to: 6729     Reply to this: 6847
Author: yuta 
Type: Update 
Category: Locking 
Subject: ETMX 1064 trans camera 

[Jenne, Yuta]

We made ETMXT camera working.
We connected the camera to video mux, placed 10% pick off mirror in front of TRX PD, lead the beam go to ETMXT camera.
Transmission to the TRY PD was 23.8 uW, but now, it's 21.3 uW (2.3 uW goes to the camera).
So, we changed C1:LSC-TRX_GAIN from -0.00181818 to -0.00203158 (=-0.00181818*23.8/21.3).

There is a channel for power normalization, C1:LSC-TRX_POW_NORM, but is 1 and it looks like we are using this gain for the normalization. Situation of TRY is the same as TRX.

ELOG V3.1.3-